1. A new pub has opened in town and you’re considering it as a potential new spot for cheap wings and beer. Before going, you chat with a group of 12 friends who have already been. You obtain the following information:

Three of them complained about the quality of the beer.

Four of them complained about the quality of the wings.

Two of them complained about the quality of the beer and wings.

Three of them had no complaints.

Answer the following questions:

a. What is the probability you will enjoy the beer?

b. What is the probability you will enjoy the wings?

c. What is the probability you will dislike either the wings, or the beer, or both?

d. What is the probability you will enjoy both the beer and the wings?


1
Expert's answer
2020-10-26T19:55:31-0400
"3+4+2+3=12"

a.


"P_1=\dfrac{12-3-2}{12}=\dfrac{7}{12}"

b.


"P_2=\dfrac{12-4-2}{12}=\dfrac{1}{2}"

c.


"P_3=\dfrac{3+4+2}{12}=\dfrac{3}{4}"

d.


"P_4=\dfrac{3}{12}=\dfrac{1}{4}"